Find the future value. (Use the Table 12.1.) (Do not round intermediate calculations. Round the "FV factor" to 4 decimal places and final answer to the nearest cent.)


Present value: $7,692.96

Length of time: 7 times

Rate: 12%

Compounded: Quarterly

Period used: 28

Rate used: 3%


FV factor used: ______________ ?


Future value_________________ ?

1
Expert's answer
2021-03-07T17:11:19-0500

The future value (FV) of an investment of present value (PV) dollars earning interest at an annual rate of r compounded m times per year for a period of t years is:

"FV=PV(1+r\/m)^{mt}"

or

"FV=PV(1+i)^{n}"

where i = r/m is the interest per compounding period and n = mt is the number of compounding periods.

We have:

"i=12/4=3\ \%" , "n=7"

Then:

FV factor used:

"(1+i)^n=1.03^7=1.2299"

Future value:

"FV=7692.96\cdot1.2299=\$\ 9461.37"